Determine the total luminous flux emitted by the source if the illumination is



15 lx at the distance 2 m from that source. The luminous intensity is equal in all directions.

1
Expert's answer
2022-01-06T10:49:18-0500


source if the illumination=15 lx

r=2 m

"I=\frac{E}{A}"

Intensity

"I=\frac{15}{4\pi r^2}"

Put value

"I=\frac{15}{4\times3.14\times4}=0.2985lx/m^2"